Please answer 2 and 3

hich of the following situations describes data from two separate populations? ( point) the grade point averages of 20 students before participating in a tutonng program compared with their grade po t geages at er pate a g in the gra O the salaries of 50 engineers in Houston compared with the salaries of so engineers in San Diego the scores of 500 eighth-grade students on a reading test in September compared to their scores in June the percentage of successful goals for 30 athletes before participating in a soccer camp and after participating in the soccer camp A school is impiementing an SAT preparation program. To study the program's efflectiveness, the school i point looks at participants' SAT scores before starting the program and after completing the program Out of 100 students sampled, 80 increased their scores, 10 kept their scores the same, and 1O decreased their scoresHow could this situation be simulated using a random digit table? Assign the digit 0 to students who increased their scores, 1 to students whose scores remained the same, and 2 to students whose scores decreased use the one-digt numbers, ignoring any values greater than 2. Assign digts o, 1,2,3,45,6, and 7 to students who i ncreased their scores. Assign s to students who kept their scoces the same and9 to students whose scores decressed. Use the random digit table to generate one-digit numbers. Assign the number 80 to students who increased ther scores, 10 to students whose scores temained the same aned 10 to stuldlents whone scores decieased. Use the random digt able generate two-digit numbers, gmoring any values besides s0 and 10. Assign digits 1,2,3,4,5,6,7,and 8 to students who decre digit table to generate one-digit numbers eased their scoresAssign9 to students wutio kept their scones thesame and 0 to students whose soores in
hich of the following situations describes data from two separate populations? (4 point) Othe grade point averages of 20 students before participating in a tutoring program compared with their grade point averages after participating in the program O the salaries of 50 engineers in Houston compared with the salaries of 50 engineers in San Diego the scores of 500 elghth-grade students on a reading test in September compared to their scores in June the percentage of successful goals for 30 athletes before participating in a soccer camp and after participating in the soccer camp 3. A school is implementing an SAT preparation program. To study the program's effectiveness, the school (2 point looks at particpants' SAT scores before starting the program and after completing the program. Out of 100 students sampled, 80 increased their scores, 10 kept their scores the same, and 10 decreased their scores. How could this situation be simulated using a random digit table? Ο Assi se ne a don a ate he same und 2 to se o ents whose scores ec eased. t a spe the d t 0 to students who increased their scores 1 to students whose scores rema ne one-digit numbers, ignoring any values greater than 2. Assign digits 0 1 2 3 4 5 6 and 7 to students who increased their scores. Assign e to students who kept their scores the same and 9 to n de es whose scores decreased. Use the random digit table to generate one-digit numbers Assign the number 80 to students who increased their scores, 10 to students whose scones remained the same,and 30 to studlents whose scores decreased, Use the random digit table to generate two-digit numbers, ignoring any valces besides 80 and 10 Assign digits 1,2,3,4,5,6,7,and s to students who decreased their scores.Assign 9 to students who kept their scones the same and o to shudents whose scores increased. Use the randorm digit table to generate one-digit numbers
Answer #1

2.
The correct answer is,
the salaries of 50 engineers in Houston compared with the salaries of 50 engineers in San Diego

Explanation - Here there ate two seperate populations - 1. engineers in Houston 2. engineers in San Diego
In all other cases, sample from same population were used in the treatments.

3.
Proportion of students who increased their scores = 80 / 100 = 80%
Proportion of students whose scores were same = 80 / 100 = 10%
Proportion of students who decreased their scores = 10 / 100 = 10%
If we generate random numbers 0-9, the total numbers to be allocated to students who increased their scores = 0.8 * 10 = 8
Total numbers to be allocated to students whose scores were same = 0.1 * 10 = 1
Total numbers to be allocated to students who decreased their scores = 0.1 * 10 = 1
So, the answer is,
Assign digits 0,1,2,3,4,5,6 and 7 to students who increased their scores. Assign 8 to students who kept their score same and 9 to students whose scores decreased. Use the random digit table to generate one-digit numbers.